OpenStudy (khally92):

I want to use proper set notation to describe a domain and my domain happens to be y^2-x^2 so in proper set notation which one is more accurate D{(x y)|y^2-x^2>0} or D{(x y)|y=-x, y=x} Thank you.

OpenStudy (khally92):

@inkyvoyd

OpenStudy (dayakar):

i think your question is incomplete, define x and y which type of numbers

OpenStudy (khally92):

\[\cos[1/\sqrt{X^2+Y^2}]\]

OpenStudy (khally92):

SORRY ITS Y^2-X^2

OpenStudy (khally92):

I don't know what i was thinking

OpenStudy (khally92):

\[\cos[1/(\sqrt{y^2-x^2})] find the domain \in proper set notation.\]

OpenStudy (khally92):

@Astrophysics

OpenStudy (khally92):

\[y^2-x^2>o\]

OpenStudy (khally92):

\[y^2=x^2\]

OpenStudy (astrophysics):

Wait what?

OpenStudy (astrophysics):

Your domain is just all real numbers then

OpenStudy (astrophysics):

y^2-x^2

OpenStudy (khally92):

\[y=x and y=-x\]

OpenStudy (astrophysics):

Oh is it cos(all that stuff)

OpenStudy (astrophysics):

\[\cos \left( \frac{ 1 }{ \sqrt{x^2-y^2} } \right)\]

OpenStudy (khally92):

exactly

OpenStudy (khally92):

y=x and y=-x ? for the domain

OpenStudy (khally92):

How can i write that in proper set notation?

OpenStudy (astrophysics):

 you wanted it like this right

OpenStudy (khally92):

how did wolfram get x^2>=y^2

OpenStudy (astrophysics):

I would've thought it was > 0 as well but looking at the alternate forms it does not seem to be the case

OpenStudy (astrophysics):

Yay jim

jimthompson5910 (jim_thompson5910):

Set the denominator equal to 0 and solve for y \[\Large \sqrt{y^2 - x^2} = 0\] \[\Large (\sqrt{y^2 - x^2})^2 = 0^2\] \[\Large y^2 - x^2 = 0\] \[\Large y^2 - x^2+x^2 = 0+x^2\] \[\Large y^2 = x^2\] \[\Large \sqrt{y^2} = \sqrt{x^2}\] \[\Large |y| = |x|\] \[\Large y = |x|\] \[\Large y = \pm x\] So if y = x or y = -x, then the denominator is equal to 0. To avoid division by zero errors, we say that y cannot equal x and y cannot equal -x. The domain is drawn from the set of real numbers R. In terms of ordered pairs, we're drawing from the set R x R = R^2. So if you use set notation, the domain D would look like this \[\Large D = \left\{(x,y) | (x,y) \in \mathbb{R}^2 \ \ , \ \ y \ne \pm x\right\}\] D is the set of ordered pairs (x,y) allowed to be plugged into the function.

jimthompson5910 (jim_thompson5910):

Sorry I'm not thinking. The radicand must also be positive, so y^2 - x^2 > 0 y^2 > x^2 |y| > |x| y > |x| or y < -|x| So the better domain is \[\Large D = \left\{(x,y) | (x,y) \in \mathbb{R}^2 \ \ , \ \ y > |x| \text{ or } y < -|x|\right\}\]

OpenStudy (astrophysics):

Haha was going to say

OpenStudy (khally92):

Perfect..And I believe the graph looks something like this